OpenStudy (anonymous):

f(x)=x^6x find f'(x) by using antilogarithm differentiation

OpenStudy (anonymous):

logarithmic not antilogarithm.. missed typed

OpenStudy (anonymous):

you have a choice, although the work is identical you can write \[x^{6x}=e^{x\ln(6x)}\] and then take the derivative using the chain rule, or you can take the log, get \[x\ln(6x)\] then take the derivative of that, and once done multiply by the original function

OpenStudy (anonymous):

all the work is the same however, you have to find the derivative of \[x\ln(6x)\] no matter which method you choose

OpenStudy (jhannybean):

Where does the x infront of ln(6x) come from? and why is that there? @satellite73

OpenStudy (anonymous):

so it would be x (6/6x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Jhannybean what is the definition of \(x^{6x}\) or for that matter how would you define \[4^{\sqrt2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

for \(x\ln(6x)\) you need the product rule

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(x)=x, f'(x)=1, g(x)=\ln(6x), g'(x)=\frac{1}{6}\]

OpenStudy (anonymous):

oops i meant \(g'(x)=\frac{1}{x}\)

OpenStudy (anonymous):

@Jhannybean i see i made a mistake, good catch

OpenStudy (agent0smith):

ln y = 6x lnx \[\huge y = e^{6x \ln x}\]

OpenStudy (anonymous):

\[x^{6x}=e^{6x\ln(x)}\]

OpenStudy (jhannybean):

Yeah, the 6 disappeared.

OpenStudy (anonymous):

yeah it was wrong, you are right also @agent0smith caught me too

OpenStudy (jhannybean):

Great explanation you guys! :D

OpenStudy (anonymous):

can't get away with anything here ok since i messed up, lets do it right you need the derivative of \(6x\ln(x)\) which still requires the product rule this time is is \[6\ln(x)+6x\times \frac{1}{x}=6\ln(x)+6\]

OpenStudy (anonymous):

i got it Thanks guys

OpenStudy (anonymous):

how do u find the derivated for 4^(x)log[7](x) @satellite73

OpenStudy (anonymous):

7 is the base of log

OpenStudy (anonymous):

the derivative of \(b^x\) is \(b^x\times \ln(b)\)

OpenStudy (anonymous):

and \[\log_b(x)=\frac{\ln(x)}{\ln(b)}\] by the all powerful change of base formula (which says all logs are the same) and so \[\frac{d}{dx}[\log_b(x)]=\frac{1}{x\ln(b)}\]

OpenStudy (anonymous):

i got the derivative, but do i need to use the product rule or not?

OpenStudy (jhannybean):

If you got the derivative why do you still need to use the product rule?

OpenStudy (anonymous):

so would the answer be 4^xln(4)(1/xln(7)) or 4^xln(4)(log_7(x))+4^(x)(1/xln(7))

OpenStudy (anonymous):

nvm.. Thanks for yall's help

OpenStudy (jhannybean):

No you're overcomplicating it :)

OpenStudy (anonymous):

lol